1. Use the following information on Cr to determine the amount of heat required to convert...

1.

Use the following information on Cr to determine the amount of heat required to convert 186.3 g of solid Cr at 1760°C into liquid Cr at 2060°C.

melting point = 1860°C; boiling point = 2672°C

ΔHfus = 20.5 kJ/mol; ΔHvap = 339 kJ/mol;

c(solid) = 44.8 J/g°C; c(liquid) = 0.94 J/g°C

Enter your answer in units of kJ to three significant figures.

2.

Approximately how many ice cubes must melt to cool 700 milliliters of water from 29°C to 0°C? Assume that each ice cube contains 1 mole of H2O and is initially at 0°C.

∆H(fusion) = 6.02 kJ/mol; ∆H(vaporization) = 40.7 kJ/mol

c(solid) = 2.09 J/g°C; c(liquid) = 4.18 J/g°C; c(gas) = 1.97 J/g°C

Enter your answer numerically.
Answer #1

Molar mass of Cr = 52 g/mol.

Moles of Cr = mass/molar mass

Heat required to increase temperature 1760 to 18600c.

H1 = moles of Cr × specific heat of solid Cr × temperature change

Or, H1 = (186.3/52) × 44.8 × (1860 -1760)

Or H1 = (3.58 × 44.8 × 100) = 16038 J = 16.038 KJ.

Heat required to melt Cr

H2 = moles of Cr × heat of fusion of Cr

Or, H2 = 3.58 × 20.5 = 73.39 KJ.

Heat required to increase the temperature 20600 c from 18600c

H3 = 3.58 × specific heat of liquid Cr × temperature change

Or, H3 = 3.58× 0.94 × (2060 - 1860)

Or, H3 = 673 J = 0.673 KJ.

Now total heat needed

= H1 +H2 + H3

= 16.038 + 73.39 + 0.673

= 90.1 KJ. (3 significant figure)

2.

Let density of water = 1 g/mL

So, mass of water = 700 mL × 1 (g/mL)

Heat released by water = heat needed to melt ice.

Heat released to decrease the temperature of water from 29 to 0oc.

q= m×C+liquid) × T

Or, q = 700 × 4.18 × (29-0)

= 84854 J

= 84.854 KJ

Now

Heat needed to melt 1 mole ice

= Moles × heat of fusion per mol

= 6.02 KJ.

So, number of ice cube must melt to cool the water

= (84.854/6.02)

= 14.09

14.

.